OpenStudy (anonymous):

(logs) find the solution:10^(x+3)=6^(x)

OpenStudy (anonymous):

start with \[(x+3)\ln(10)=x\ln(6)\] and then do some algebra to solve for x

OpenStudy (anonymous):

i need help with doing the rest, i know how to do that but then i get lost

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[\ln(10)x+3\ln(10)=\ln(6)x\] \[\ln(10)x-\ln(6)x=-3\ln(10)\] \[x(\ln(10)-\ln(6))=-3\ln(10)\] \[x=\frac{-3\ln(10)}{\ln(10)-\ln(6)}\]

OpenStudy (anonymous):

don't forget that \[\ln(10)\] and \[\ln(6)\] are just some numbers, treat them as constants

OpenStudy (anonymous):

i really dont get it:/

OpenStudy (anonymous):

can you explain it to me please